OpenStudy (anonymous):

a coin is loaded so that the probability of getting a heads is 3/4 if the coin is flipped twice what is probability of getting head twice

OpenStudy (anonymous):

\(\frac{3}{4}\times \frac{3}{4}\)

OpenStudy (anonymous):

cross multiply?

OpenStudy (anonymous):

there is no such thing as "cross multiply" no matter what those math teachers told you it does not exist multiply means multiply for fractions that means straight across

OpenStudy (anonymous):

\[\frac{a}{b}\times \frac{c}{d}=\frac{a\times c}{b\times d}\] \[\frac{3}{4}\times \frac{3}{4}=\frac{3\times 3}{4\times 4}\]

OpenStudy (anonymous):

\[3/4*3/4=(3*3)/(4*4)=9/16\]

OpenStudy (anonymous):

yes 9/16

OpenStudy (anonymous):

that is it, yes

OpenStudy (anonymous):

i need to simplify that right

OpenStudy (anonymous):

i am going to sneak out one day and smack every math teacher that ever said "cross multiply" "cross cancel" and "simplify" there is no such thing you can in some cases "reduce to lowest terms" but not in this case because there are no common factors of \(9\) and \(16\)

OpenStudy (anonymous):

that is not a answer choice

OpenStudy (anonymous):

It is, however, the correct answer. Is 0.5625 an answer choice?

OpenStudy (anonymous):

that's too bad what are the choices?

OpenStudy (anonymous):

wait yes it is! sorry thanks!!

OpenStudy (anonymous):

Good luck! :)